The average price of seven ABC dishwashers was $845, and the average price of nine XYZ dishwashers was $815. The standard deviations were $9 and $19, respectively. At alpha = 0.05, can one conclude that the ABC dishwasher cost more?

Ho:μ ABC = μ XYZ

H1: μ ABC > μ XYZ

a. Calculate the Test Statistic.

b. Find the P-value.
Answer #1

Below are the null and alternative Hypothesis,
Null Hypothesis, H0: μ1 = μ2
Alternative Hypothesis, Ha: μ1 > μ2

Pooled Variance
sp = sqrt(s1^2/n1 + s2^2/n2)
sp = sqrt(81/7 + 361/9)
sp = 7.1891

Test statistic,
z = (x1bar - x2bar)/sp
z = (845 - 815)/7.1891
z = 4.17

P-value Approach
P-value = 0

As P-value < 0.05, reject the null hypothesis.